Whats the probability that if you draw 2 marbles without replacement out of a bag containing 1 green, 2 yellow, and 3 red, that both of them are red?
Answer by stanbon(75887)

Whats the probability that if you draw 2 marbles without replacement out of a bag containing 1 green, 2 yellow, and 3 red, that both of them are red?
----
# of ways to succeed: 3C2 = 3
# of random pairs: 6C2 = 15
---
P(2 red) = 3/15 = 1/5
=========================
Cheers,
Stan H.
===================